OpenStudy (anonymous):

10d-6=4d-15-3d ((show work))

OpenStudy (alekos):

first you need to reduce the RHS

OpenStudy (imtiaz7):

10d-6=d-15 10d-d=-15+6 9d=-9 d=-9/9 d=-1